Suppose that you deposit $1,000 in a savings account at Wells Fargo and plan to leave your principal and any interest in the account for four months. The interest rate in the first month is 3%; in the second month, 4%; in the third month, 2%; and in the fourth month, 1%. What is the future value of your account at the end of the holding period? Do not round at intermediate steps in your calculation. Round your final answer to the nearest penny. Do not type the $ symbol.
Answer #1

Future Value = PV *(1+3%)*(1+4%)*(1+2%)*(1+1%) = 1000*1.03*1.04*1.02*1.01 = 1103.55

Assuming second month rate is 3%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well